OpenStudy (anonymous):

The width of a rectangle is one half its length. the perimeter of the rectangle is 54 cm. what are the width and length of the rectangle.

OpenStudy (anonymous):

please explain how u got ur answer

OpenStudy (anonymous):

If the width is w, what is the length (in terms of w)?

OpenStudy (anonymous):

w = l/2 p = 2w + 2l = 54 =l + 2l 3l = 54 l = 18 w = 9

OpenStudy (anonymous):

can you explain blah

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Answer the question I asked earlier.